# HOFT: Householder Orthogonal Fine-tuning

### v1

@@ 1 Introduction
Nowadays, fine-tuning foundation models for downstream tasks [17] is the standard approach to model adaptation these days thanks to their knowledge across many domains. By tuning far fewer parameters than full fine-tuning using parameter-efficient fine-tuning techniques (PEFT), the model is able to learn the key aspects of a task and perform comparably or even better than full fine-tuning [18]. This fact makes PEFT methods a particularly well-suited approach for efficiently adapting these models. The employment of PEFT methods has enabled the adaptation of large foundation models without the necessity of compute-intensive hardware infrastructure, making adaptation accessible to a broader user community.
@@
The most popular PEFT methods are based on low-rank approximations, including Low-Rank Adaptation (LoRA) [19] and Weight-Decomposed Low-Rank Adaptation (DoRA) [27]. These methods work under the assumption that the learnable parameters must reside in a lower intrinsic dimension [2]. Alternatively, there are methods proposing the use of orthogonal matrices for adaptation, such as Orthogonal Fine-tuning (OFT) [36] and Orthogonal Butterfly (BOFT) [29]. These methods hypothesize that a good fine-tuned model should have a minimal difference in hyperspherical energy compared to the pre-trained model [26, 28]. In brief, the assumption made is that orthogonality is required to learn new features while keeping pre-trained information [36, 53]. Whilst the performance of these techniques has been demonstrated, their runtime and memory footprint make them a less preferable option for use in real-world applications. A recent approach to balance low-rank and orthogonal methods is Householder Reflection Adaption (HRA) [53], which constrains orthogonality through the incorporation of a term within the loss function. With the employment of an additional weight $\lambda$ for the orthogonality regularizer, HRA aims to construct the chained matrix product of Householder transformations [14].
@@
Orthogonal fine-tuning methods generally result in the construction of a single orthogonal matrix for adaptation purposes. This work demonstrates that two orthogonal matrices are required in order to ensure full expressivity in orthogonal fine-tuning methods. This leads us to propose Householder Orthogonal Fine-tuning (HOFT): a novel orthogonal fine-tuning technique using two orthogonal matrices as directional components efficiently updated through orthogonal transformations. For efficiency, these matrices are obtained by accumulating Householder transformations via the CWY transform [25, 21] along with a fast inverse approximation. Additionally, we draw inspiration from DoRA’s analysis, which shows that fine-tuning magnitude and direction separately closely matches the learning dynamics of full fine-tuning. From this, a variant of HOFT incorporating an additional scaling transformation is proposed: Scaled Householder Orthogonal Fine-tuning (SHOFT).
@@ 2 Related Work
Methods in this family assume that effective fine-tuning updates lie on a compact, low-dimensional manifold [19, 27, 23, 55, 24, 51, 20, 46]. LoRA [19] introduces trainable low-rank adapter matrices into each Transformer layer, freezing the original weights and reducing trainable parameters by several orders of magnitude. DoRA [27] separates the fine-tuning of directional and scaling components by normalizing LoRA’s output and applying a scaling transformation. PiSSA [30] employs singular value decomposition (SVD) on pre-trained weight matrices to initialize LoRA adapters in principal subspaces, maintaining most of the original model’s expressive capacity. QLoRA [12] combines 4-bit NormalFloat (NF4) quantization with LoRA, enabling the fine-tuning of 65B-parameter models on a single 48GB GPU while preserving near full-precision quality. QA-LoRA [49] employs group-wise quantization operators to selectively compress adapter updates with minimal impact on task performance loss.
@@
Orthogonal fine-tuning methods learn distance preserving transformations in weight space, keeping geometric properties such as hyperspherical energy among neuron activations [36, 29]. Previous works show how the imposition of orthogonality constraints within deep learning architectures is conducive to enhancing performance [5, 44, 48, 13, 3]. OFT [36] employs Cayley parameterization [22] to generate orthogonal matrix blocks. Additionally, COFT [36] constrains the orthogonal matrix to be within a small neighborhood of the pre-trained matrix. BOFT [29] reduces OFT parameter footprint by factorizing orthogonal updates into butterfly structures inspired by the Cooley–Tukey FFT algorithm [6], achieving similar generalization gains with fewer trainable parameters. The employment of hybrid methods, such as HRA, enforces hyperspherical constraints on low-rank adapters to blend both paradigms via a term in the loss function controlled by a weight [53].
@@ 3.1 Orthogonal fine-tuning paradigm
As discussed in Section 2, orthogonal fine-tuning stresses the importance of preserving the hyperspherical energy of the given matrix $\mathbf{M}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}\in\mathbb{R}^{m\times n}$. Although it is clear that this can be done by adapting both singular vector matrices $\mathbf{U}$ and $\mathbf{V}$, it is common practice to keep $\mathbf{V}^{\top}$ unchanged and adapt only $\mathbf{U}$ [36, 29, 53].
@@
In terms of approximation error, pre- and post-multiplying the pre-trained matrix $\mathbf{M}$ by distance-preserving transformations exactly captures all adapted matrices $\widehat{\mathbf{M}}$ that maintain the same hyperspherical energy. However, the error incurred when applying just one orthogonal matrix leads us to a known problem, the Orthogonal Procrustes Problem [15], which has a solution if $\mathbf{\widehat{M}}$ and $\mathbf{M}$ are known matrices. In this one-transform setting, a theoretical upper bound on the Frobenius norm error is given by
@@ 3.2 CWY transform and inverse approximation
As observed in [36, 29], computing parameterized orthogonal matrices is computationally costly though it can be sped up with numerical methods. In our case, the composition of multiple Householder transformations can be cast into high-performance matrix-matrix products through the WY and CWY transforms [21, 25]. The following result from [21] allows us to construct an orthogonal matrix by accumulating householder transformations:
@@
As in the case of many orthogonal parameterization methods [14], there is a matrix inverse to be computed. This fact makes orthogonal parameterization methods non-scalable, since the inverse computation during training and the gradient update computation are resource-intensive. However, in the case of the CWY transform, the inverse can be approximated with a high degree of precision. In order to efficiently compute $\mathbf{S}^{-1}$, Neuman Series are required [14]. We can separate $\mathbf{S}=\mathbf{D}+\mathbf{A}=\mathbf{D}(\mathbf{I}+\mathbf{D}^{-1}\mathbf{A})$ where $\mathbf{D}$ is a diagonal matrix and $\mathbf{A}$ is a strictly upper triangular matrix. The inverse will be:
@@ 3.3 Householder Orthogonal Fine-tuning
Similar to HRA’s rank $r$ [53], HOFT also employs $r$ householder vectors. For both inverse approximations, the computational complexity is $\mathcal{O}\left(r^{2}(m+n)\right)$, and the matrix-vector multiplications require $\mathcal{O}\left(2mr+2nr+mn+2r^{2}\right)$. Altogether, the total time complexity of HOFT is $\mathcal{O}\left(mn+2r(m+n)+r^{2}(m+n+2)\right)\sim\mathcal{O}\left(mn+(m+n)(r^{2}+2r)\right)$. A comparison of the computational complexity of HOFT to other parameterized orthogonal-based methods is provided in Table 1.
@@
One drawback of OFT is that it requires $b$ to be large in order to achieve $\text{O}(m)$ coverage [36]. The increase of $b$ cannot be arbitrary because of the cost of inverting $b\times b$ matrices. BOFT, on the other hand, offers better coverage at the expense of higher time complexity [29]. HRA provides even better coverage than the two previous methods; however, its Householder transformations must be applied sequentially, and when $\lambda=\infty$, its runtime matches that of OFT [53]. By contrast, HOFT provides the same coverage as HRA, and because most of its computations can be parallelized, it achieves greater speedup and represents an attractive alternative.
Although LoRA and DoRA can be randomly initialized, OFT and BOFT cannot due to the necessity of preserving orthogonality; Cayley’s parameterization [14] needs skew-symmetric matrix $\mathbf{R}=\boldsymbol{0}$ to ensure that the orthogonal parameterized matrix is $\mathbf{Q}=\mathbf{I}$. In general, orthogonal PEFT methods cannot be randomly initialized. However, HOFT and SHOFT can be randomly initialized by considering consecutive pairs of equal vectors $\mathbf{u}_{i}$. Since they express the same reflection, we can place Householder vectors in the form $\mathbf{U}=(\mathbf{u}_{1}\;|\;\mathbf{u}_{1}\;|\;\cdots\;|\;\mathbf{u}_{k}\;|\;\mathbf{u}_{k})$, which yields the identity matrix:
@@ 3.4 Scaled Householder Orthogonal Fine-tuning
The use of scaling transformations in orthogonal fine-tuning methods has been studied in [36] as a way to improve their performance. Drawing also inspiration from DoRA’s weight decomposition analysis [27], we propose a variant of HOFT that employs a scaling transformation: Scaled Householder Orthogonal Fine-tuning (SHOFT). As observed in Section 3.1, placing the scaling transformation near the singular value matrix will be interesting from a SVD perspective. Since scaling is performed between two distance preserving transformations, the effect of $\mathbf{m}$ in the singular values of $\mathbf{M}$ is closely controlled. Thus, SHOFT formulation will be as follows
@@
where $\mathbf{Q}_{\mathbf{U}},\mathbf{Q}_{\mathbf{V}}$ and $\mathbf{m}$ are formed by trainable parameters. It seems more intuitive to be able to redirect with $\mathbf{Q}_{\mathbf{V}}$, transform with $\mathbf{M}$, then scale with $\mathbf{m}$ and finally redirect with $\mathbf{Q}_{\mathbf{U}}$. SHOFT is more flexible since it is no longer constrained to keep the same hyperspherical energy. All elements of vector $\mathbf{m}$ are initialized to one. As observed in other PEFT methods [27, 36], the increase on the amount of trainable parameters due to adding a magnitude vector $\mathbf{m}\in\mathbb{R}^{m}$ is marginal.
@@ 4.1 Commonsense reasoning
For measuring commonsense reasoning performance, we compare HOFT and SHOFT with DoRA and LoRA across eight standard commonsense reasoning benchmarks: BoolQ [8], PIQA [4], SIQA [43], HellaSwag [54], WinoGrande [42], ARC-e [9], ARC-c [9] and OBQA [31]. Following DoRA [27], the training splits of all eight tasks are merged into a single training set, and then each model is evaluated separately on the original test set of each task. The models employed are LLaMA3.1-8B [16], Qwen2.5-7B [50], Phi4-14B [1], and Qwen2.5-14B [50]. We initialize DoRA [27] and LoRA [19] using PiSSA [30]. We set $r=16$ for all PEFT methods and train the models for two epochs.
The results of each individual task along with the average task accuracy per model and PEFT method are shown in Table 2, where it can be seen that HOFT and SHOFT generally achieve higher scores than LoRA and DoRA across most models, with SHOFT performing comparably to DoRA for Qwen2.5-7B. Moreover, both HOFT and SHOFT continue to deliver strong results as model size grows, demonstrating solid performance on both Phi4-14B and Qwen2.5-14B. In particular, HOFT and SHOFT attain the highest scores on nearly every task, matching LoRA and DoRA only on PIQA and ARC-e. This underscores their robustness and efficiency when trained on datasets containing multiple domains.

### v2

@@ 1 Introduction
Nowadays, fine-tuning foundation models for downstream tasks (He et al., 2021) is the standard approach to model adaptation these days thanks to their knowledge across many domains. By tuning far fewer parameters than full fine-tuning using parameter-efficient fine-tuning techniques (PEFT), the model is able to learn the key aspects of a task and perform comparably or even better than full fine-tuning (Houlsby et al., 2019). This fact makes PEFT methods a particularly well-suited approach for efficiently adapting these models. The employment of PEFT methods has enabled the adaptation of large foundation models without the necessity of compute-intensive hardware infrastructure, making adaptation accessible to a broader user community.
@@
The most popular PEFT methods are based on low-rank approximations, including Low-Rank Adaptation (LoRA) (Hu et al., 2022) and Weight-Decomposed Low-Rank Adaptation (DoRA) (Liu et al., 2024a). These methods work under the assumption that the learnable parameters must reside in a lower intrinsic dimension (Aghajanyan et al., 2021). Alternatively, there are methods proposing the use of orthogonal matrices for adaptation, such as Orthogonal Fine-tuning (OFT) (Qiu et al., 2023) and Orthogonal Butterfly (BOFT) (Liu et al., 2024b). These methods hypothesize that a good fine-tuned model should have a minimal difference in hyperspherical energy compared to the pre-trained model (Lin et al., 2020; Liu et al., 2018). In brief, the assumption made is that orthogonality is required to learn new features while keeping pre-trained information (Qiu et al., 2023; Yuan et al., 2024). Whilst the performance of these techniques has been demonstrated, their runtime and memory footprint make them a less preferable option for use in real-world applications. A recent approach to balance low-rank and orthogonal methods is Householder Reflection Adaption (HRA) (Yuan et al., 2024), which constrains orthogonality through the incorporation of a term within the loss function. With the employment of an additional weight $\lambda$ for the orthogonality regularizer, HRA aims to construct the chained matrix product of Householder transformations (Golub and Van Loan, 2013).
@@
Orthogonal fine-tuning methods generally result in the construction of a single orthogonal matrix for adaptation purposes. This work demonstrates that two orthogonal matrices are required in order to ensure full expressivity in orthogonal fine-tuning methods. This leads us to propose Householder Orthogonal Fine-tuning (HOFT): a novel orthogonal fine-tuning technique using two orthogonal matrices as directional components efficiently updated through orthogonal transformations. For efficiency, these matrices are obtained by accumulating Householder transformations via the CWY transform (Likhosherstov et al., 2021; Joffrain et al., 2006) along with a fast inverse approximation. Additionally, we draw inspiration from DoRA’s analysis, which shows that fine-tuning magnitude and direction separately closely matches the learning dynamics of full fine-tuning. From this, a variant of HOFT incorporating an additional scaling transformation is proposed: Scaled Householder Orthogonal Fine-tuning (SHOFT).
@@ 2 Related Work
Methods in this family assume that effective fine-tuning updates lie on a compact, low-dimensional manifold (Hu et al., 2022; Liu et al., 2024a; Kopiczko et al., 2024; Zhang et al., 2023; Li et al., 2024; Yeh et al., 2023; Hyeon-Woo et al., 2022; Valipour et al., 2023). LoRA (Hu et al., 2022) introduces trainable low-rank adapter matrices into each Transformer layer, freezing the original weights and reducing trainable parameters by several orders of magnitude. DoRA (Liu et al., 2024a) separates the fine-tuning of directional and scaling components by normalizing LoRA’s output and applying a scaling transformation. PiSSA (Meng et al., 2024) employs singular value decomposition (SVD) on pre-trained weight matrices to initialize LoRA adapters in principal subspaces, maintaining most of the original model’s expressive capacity. QLoRA (Dettmers et al., 2024) combines 4-bit NormalFloat (NF4) quantization with LoRA, enabling the fine-tuning of 65B-parameter models on a single 48GB GPU while preserving near full-precision quality. QA-LoRA (Xu et al., 2024) employs group-wise quantization operators to selectively compress adapter updates with minimal impact on task performance loss.
@@
Orthogonal fine-tuning methods learn distance preserving transformations in weight space, keeping geometric properties such as hyperspherical energy among neuron activations (Qiu et al., 2023; Liu et al., 2024b). Previous works show how the imposition of orthogonality constraints within deep learning architectures is conducive to enhancing performance (Boissin et al., 2025; Sui et al., 2024; Wu et al., 2023; Fei et al., 2022; Biegun et al., 2024). OFT (Qiu et al., 2023) employs Cayley parameterization (Khare and Stewart, 2019) to generate orthogonal matrix blocks. Additionally, COFT (Qiu et al., 2023) constrains the orthogonal matrix to be within a small neighborhood of the pre-trained matrix. BOFT (Liu et al., 2024b) reduces OFT parameter footprint by factorizing orthogonal updates into butterfly structures inspired by the Cooley–Tukey FFT algorithm (Brigham, 1974), achieving similar generalization gains with fewer trainable parameters. The employment of hybrid methods, such as HRA, enforces hyperspherical constraints on low-rank adapters to blend both paradigms via a term in the loss function controlled by a weight (Yuan et al., 2024).
@@ 3.1 Orthogonal fine-tuning paradigm
As discussed in Section 2, orthogonal fine-tuning stresses the importance of preserving the hyperspherical energy of the given matrix $\mathbf{M}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}\in\mathbb{R}^{m\times n}$. Although it is clear that this can be done by adapting both singular vector matrices $\mathbf{U}$ and $\mathbf{V}$, it is common practice to keep $\mathbf{V}^{\top}$ unchanged and adapt only $\mathbf{U}$ (Qiu et al., 2023; Liu et al., 2024b; Yuan et al., 2024).
@@
In terms of approximation error, pre- and post-multiplying the pre-trained matrix $\mathbf{M}$ by distance-preserving transformations exactly captures all adapted matrices $\widehat{\mathbf{M}}$ that maintain the same hyperspherical energy. However, the error incurred when applying just one orthogonal matrix leads us to a known problem, the Orthogonal Procrustes Problem (Gower and Dijksterhuis, 2004), which has a solution if $\mathbf{\widehat{M}}$ and $\mathbf{M}$ are known matrices. In this one-transform setting, a theoretical upper bound on the Frobenius norm error is given by
@@ 3.2 CWY transform and inverse approximation
As observed in (Qiu et al., 2023; Liu et al., 2024b), computing parameterized orthogonal matrices is computationally costly though it can be sped up with numerical methods. In our case, the composition of multiple Householder transformations can be cast into high-performance matrix-matrix products through the WY and CWY transforms (Joffrain et al., 2006; Likhosherstov et al., 2021). The following result from (Joffrain et al., 2006) allows us to construct an orthogonal matrix by accumulating householder transformations:
@@
As in the case of many orthogonal parameterization methods (Golub and Van Loan, 2013), there is a matrix inverse to be computed. This fact makes orthogonal parameterization methods non-scalable, since the inverse computation during training and the gradient update computation are resource-intensive. However, in the case of the CWY transform, the inverse can be approximated with a high degree of precision. In order to efficiently compute $\mathbf{S}^{-1}$, Neuman Series are required (Golub and Van Loan, 2013). We can separate $\mathbf{S}=\mathbf{D}+\mathbf{A}=\mathbf{D}(\mathbf{I}+\mathbf{D}^{-1}\mathbf{A})$ where $\mathbf{D}$ is a diagonal matrix and $\mathbf{A}$ is a strictly upper triangular matrix. The inverse will be:
@@ 3.3 Householder Orthogonal Fine-tuning
Similar to HRA’s rank $r$ (Yuan et al., 2024), HOFT also employs $r$ householder vectors. For both inverse approximations, the computational complexity is $\mathcal{O}\left(r^{2}(m+n)\right)$, and the matrix-vector multiplications require $\mathcal{O}\left(2mr+2nr+mn+2r^{2}\right)$. Altogether, the total time complexity of HOFT is $\mathcal{O}\left(mn+2r(m+n)+r^{2}(m+n+2)\right)\sim\mathcal{O}\left(mn+(m+n)(r^{2}+2r)\right)$. A comparison of the computational complexity of HOFT with other parameterized orthogonal-based methods is presented in Table 1. The table also reports the number of parameters required by each method, together with their corresponding orthogonal coverage. We say that a method covers the orthogonal group $O(m)$ if its parameterization function is surjective, that is, it can express all $O(m)$ matrices.
@@
One drawback of OFT is that it requires the block size $b$ to be large in order to achieve $\text{O}(m)$ coverage (Qiu et al., 2023). As $b$ increases, the block diagonal orthogonal matrix is closer to a full orthogonal matrix. The increase of $b$ cannot be arbitrary because of the cost of inverting $b\times b$ matrices. BOFT, on the other hand, offers better coverage at the expense of higher time complexity (Liu et al., 2024b). HRA provides even better coverage than the two previous methods; however, its Householder transformations must be applied sequentially, and when $\lambda=\infty$, its runtime matches that of OFT (Yuan et al., 2024). By contrast, HOFT provides the same coverage as HRA, and because most of its computations can be parallelized, it achieves greater speedup and represents an attractive alternative.
Although LoRA and DoRA can be randomly initialized, OFT and BOFT cannot due to the necessity of preserving orthogonality; Cayley’s parameterization (Golub and Van Loan, 2013) needs skew-symmetric matrix $\mathbf{R}=\boldsymbol{0}$ to ensure that the orthogonal parameterized matrix is $\mathbf{Q}=\mathbf{I}$. In general, orthogonal PEFT methods cannot be randomly initialized. However, HOFT and SHOFT can be randomly initialized by considering consecutive pairs of equal vectors $\mathbf{u}_{i}$. Since they express the same reflection, we can place Householder vectors in the form $\mathbf{U}=(\mathbf{u}_{1}\;|\;\mathbf{u}_{1}\;|\;\cdots\;|\;\mathbf{u}_{k}\;|\;\mathbf{u}_{k})$, which yields the identity matrix:
@@ 3.4 Scaled Householder Orthogonal Fine-tuning
The use of scaling transformations in orthogonal fine-tuning methods has been studied in (Qiu et al., 2023) as a way to improve their performance. Drawing also inspiration from DoRA’s weight decomposition analysis (Liu et al., 2024a), we propose a variant of HOFT that employs a scaling transformation: Scaled Householder Orthogonal Fine-tuning (SHOFT). As observed in Section 3.1, placing the scaling transformation near the singular value matrix will be interesting from a SVD perspective. Since scaling is performed between two distance preserving transformations, the effect of $\mathbf{m}$ in the singular values of $\mathbf{M}$ is closely controlled. Thus, SHOFT formulation will be as follows
@@
where $\mathbf{Q}_{\mathbf{U}},\mathbf{Q}_{\mathbf{V}}$ and $\mathbf{m}$ are formed by trainable parameters. It seems more intuitive to be able to redirect with $\mathbf{Q}_{\mathbf{V}}$, transform with $\mathbf{M}$, then scale with $\mathbf{m}$ and finally redirect with $\mathbf{Q}_{\mathbf{U}}$. SHOFT is more flexible since it is no longer constrained to keep the same hyperspherical energy. All elements of vector $\mathbf{m}$ are initialized to one. As observed in other PEFT methods (Liu et al., 2024a; Qiu et al., 2023), the increase on the amount of trainable parameters due to adding a magnitude vector $\mathbf{m}\in\mathbb{R}^{m}$ is marginal.
@@ 4.1 Commonsense reasoning
For measuring commonsense reasoning performance, we compare HOFT and SHOFT with DoRA, LoRA and OFT across eight standard commonsense reasoning benchmarks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e (Clark et al., 2018), ARC-c (Clark et al., 2018) and OBQA (Mihaylov et al., 2018). Following DoRA (Liu et al., 2024a), the training splits of all eight tasks are merged into a single training set, and then each model is evaluated separately on the original test set of each task. The models employed are LLaMA3.1-8B (Grattafiori et al., 2024), Qwen2.5-7B (Yang et al., 2024), Phi4-14B (Abdin et al., 2024), and Qwen2.5-14B (Yang et al., 2024). We initialize DoRA (Liu et al., 2024a) and LoRA (Hu et al., 2022) using PiSSA (Meng et al., 2024). We set $r=16$ for all PEFT methods and train the models for two epochs.
@@
The results of each individual task along with the average task accuracy per model and PEFT method are shown in Table 2, where it can be seen that HOFT and SHOFT generally achieve higher scores than LoRA, DoRA and OFT across most models, with SHOFT performing comparably to DoRA for Qwen2.5-7B. Moreover, both HOFT and SHOFT continue to deliver strong results as model size grows, demonstrating solid performance on both Phi4-14B and Qwen2.5-14B. In particular, HOFT and SHOFT attain the highest scores on nearly every task, matching LoRA and DoRA only on PIQA and ARC-e. This underscores their robustness and efficiency when trained on datasets containing multiple domains.
Table 3: Performance comparison on $\text{X}\rightarrow\text{English}$ machine translation tasks
Model Method #Params (%) Slovene German Latvian French Average
BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET
NLLB-3.3B Baseline - 39.7 87.5 39.3 86.2 31.2 81.3 38.5 84.9 37.2 85.0
LoRA 0.42 46.8 89.2 44.5 87.7 38.2 83.9 49.7 87.8 44.8 87.2
DoRA 0.43 46.8 89.1 44.7 87.6 38.2 83.9 49.5 87.7 44.8 87.1
OFT 0.43 45.5 89.1 44.2 87.5 36.9 83.3 49.1 87.6 43.9 86.9
BOFT 0.43 45.3 89.1 44.1 87.5 36.9 83.4 49.1 87.7 43.9 86.9
HOFT 0.42 48.0 89.4 44.4 87.6 38.6 83.9 49.5 87.7 45.1 87.2
SHOFT 0.43 46.4 89.5 44.5 87.7 38.7 84.0 49.7 87.8 44.8 87.3
LLaMA2-7B 0-shot - 26.8 72.8 30.4 74.1 4.5 52.2 37.2 79.3 24.7 69.6
LoRA 0.19 39.3 84.7 41.5 86.9 15.5 66.2 47.0 87.2 35.8 81.3
DoRA 0.19 39.6 84.8 41.4 86.9 16.2 66.6 47.0 87.2 36.1 81.4
OFT 0.19 40.3 84.7 41.3 86.8 11.0 61.9 46.9 87.2 34.9 80.2
BOFT 0.19 40.4 84.9 41.2 86.8 11.0 61.6 46.9 87.2 34.9 80.1
HOFT 0.19 40.6 85.2 41.4 86.9 15.8 66.6 47.0 87.3 36.2 81.5
SHOFT 0.19 41.2 85.6 41.6 87.0 15.7 65.9 47.1 87.3 36.4 81.5
LLaMA3.1-8B 0-shot - 34.2 77.8 40.9 86.2 22.9 70.8 41.6 82.7 34.9 79.4
LoRA 0.12 36.2 84.1 42.3 87.4 32.7 80.9 46.8 85.5 39.5 84.5
DoRA 0.12 42.4 85.0 42.2 87.4 32.8 80.8 46.7 85.5 41.0 84.7
OFT 0.16 21.5 74.7 43.5 87.6 17.4 69.4 46.8 85.5 32.3 79.3
BOFT 0.16 22.1 75.9 43.6 87.6 17.1 69.2 46.8 85.5 32.4 79.6
HOFT 0.12 44.2 86.6 42.9 87.5 32.2 80.4 46.7 85.6 41.5 85.0
SHOFT 0.12 43.6 86.4 43.1 87.7 31.9 80.4 46.8 85.6 41.4 85.0